Question 1 (1 point)

Electromagnetic waves are generally

Question 1 options:

transverse waves

standing waves.

longitudinal waves.

50/50 combination of transverse and longitudinal waves.

sub section

Question 2 (1 point)

The period of a sound wave is the

Question 2 options:

decade in which it is produced.

None of these.

inverse of its frequency.

end of the wave.

sub section

Question 3 (1 point)

If the speed of sound is 1100 feet per second and a tuning fork oscillates at 440 cycles per second, what is the wavelength of the sound produced?

Question 3 options:

0.4 feet

2.5 feet

4.4 feet

11 feet

sub section

Question 4 (1 point)

On a day when the speed of sound in air is 340 m/s, a lightning bolt is observed and the resultant thunderclap is heard 5 seconds afterwards. About how far away was the lightning bolt?

Question 4 options:

68 m

1700 m

1020 m

88,000 m

sub section

Question 5 (1 point)

You hear a musician playing a series of notes of increasing frequency. As the frequency increases, the wave speed

Question 5 options:

stays the same.

decreases.

increases.
Answer #1

Answers:

Question (1) Answer : transverse waves

Question (2) Answer : inverse of its frequency.

Question (3) Answer : 2.5 feet

Since: v = \lambda f

Thus: \lambda = v / f = (1100 feet/sec) / (440 s-1) = 2.5 feet

Question (4) Answer : 1700 m

Question (5) Answer : stays the same.